OpenStudy (anonymous):

Cubic function help? x^3+x+31

OpenStudy (jdoe0001):

so... what are we doing again?

OpenStudy (haseeb96):

3x^2 +1 is the answer of it

OpenStudy (anonymous):

Me and a guy have been stuck on this problem for like an hour. He was given a horrible example for a cubic function, and is determined to figure it out. But it's weird, and missing a peice, so you can't factor by grouping. It's x^3+x+31=0. How did you find that @Haseeb96

OpenStudy (jdoe0001):

x^3+x+31 = 3x^2 +1 ?

OpenStudy (jdoe0001):

x^3+x+31=0 so find "x" ?

OpenStudy (anonymous):

How? We need an explanation.. I guess.

OpenStudy (anonymous):

How did both of you get 3x^2+1?

OpenStudy (jdoe0001):

hmm I didn't, I wonder that myself

OpenStudy (anonymous):

3x^2+1 is the derivative of that function.. Exactly what is it that you are asked to do?

OpenStudy (anonymous):

I think he's explaining it. What do you mean by derivative?

OpenStudy (jdoe0001):

something tells me isn't a calculus question

OpenStudy (anonymous):

I take it you're not in calculus yet, so that won't help you too much. haha

OpenStudy (anonymous):

I don't really know. He was given that bobo example. I don't know why he couldn't just give him a regular one. I'm in Pre-Cal Honors.

OpenStudy (jdoe0001):

This_Is_Batman derivative is a way to get a slope of a curve, calculus material though

OpenStudy (haseeb96):

f(x)=x^3+x+31 you are going to find its derivative according to the formula of Power function since,dx^n\dx = nx^n-1 df(x)\dx =d\dx (x^3+x+31) apply power function in it = 3x^2+1 +0 =3x^2+1 this is the solution of it

OpenStudy (anonymous):

Oh, no. I'm homeschooled, so I'm pretty much teaching myself calculus... :c

OpenStudy (anonymous):

@SolomonZelman Look at what Hasee said.

OpenStudy (solomonzelman):

Lmao, DUDE!!!!!!!

OpenStudy (anonymous):

That's what that first girl said, she just didn't fully explain it. -_-

OpenStudy (solomonzelman):

@Haseeb96, can you explain this, or you just googled it, and got the answer thinking that the website you clicked is answering this question, when it was answering something else?

OpenStudy (anonymous):

Yes, I would like it explained more thouroughly, too. I'm taking this class this year, and would like to be prepared. c:

OpenStudy (solomonzelman):

@Haseeb96, maybe you can explain this?

OpenStudy (solomonzelman):

\(\huge {\bbox[5pt,yellow,border:2px solid yellow]{~•~~‿~•~}}\)

OpenStudy (haseeb96):

boys , this is the chapter of differtiablity. in which we take the deriavtive of the function and get the answer....

OpenStudy (solomonzelman):

\(\huge\color{blue}{\huge {\bbox[5pt,cyan,border:2px solid purple]{~•~~‿~•~}}}\)

OpenStudy (anonymous):

I am a female. -_-

OpenStudy (solomonzelman):

I am a male, senior, turning 17.

OpenStudy (solomonzelman):

I meant mail.

OpenStudy (anonymous):

I'm a female, 16, Junior, pregnant. Who cares? :p And what do you mean you meant mail??? You spelled it right the first time..

OpenStudy (haseeb96):

i am also 17 years old....

OpenStudy (anonymous):

Well, hasee is apparently 17, asian, and going to bed. -_-

OpenStudy (anonymous):

Sorry @SolomonZelman , I really wanted to figure this out, but I have no idea how to do that derivative thing yet...

OpenStudy (solomonzelman):

Me niether.

OpenStudy (anonymous):

Before you get into doing derivates, it's important to have a good understanding of limits. Have you spent any time with those?

OpenStudy (anonymous):

derivatives**

OpenStudy (anonymous):

Nope. But I'm a year ahead in math, so I don't feel so dumb. :p Lol. I'm on FLVS, and only about a quater of the way through... Do you have a simple explanation? :D

OpenStudy (solomonzelman):

\(\Huge{\color{purple}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{cyan}{\bigstar}\color{purple}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{cyan}{\bigstar}\color{purple}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\sf\color{magenta}{~~I~~~suck~~~at~~~MATH!~\\}\\\color{white}{.}\\\\\Huge{\color{purple}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{cyan}{\bigstar}\color{purple}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{cyan}{\bigstar}\color{purple}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

OpenStudy (anonymous):

Really? -.- You have way too much fun on this website. :p

OpenStudy (solomonzelman):

\(\huge\color{yellow}{\huge {\bbox[5pt,red,border:2px solid magenta ]{•‿•}}}\)

OpenStudy (anonymous):

@Helloimjohn1234

OpenStudy (anonymous):

I'm not sure what FLVS means.. haha. Limits are pretty simple at their root, but are used throughout college math. I'm currently finishing up Calc 3 and I still use them regularly. The concepts are too long and detailed for me to type out on this page, but this website will probably be really helpful to you. Limits are the first thing you cover in Calc 1 so that's where this page starts. Do some reading!! http://tutorial.math.lamar.edu/Classes/CalcI/LimitsIntro.aspx

OpenStudy (anonymous):

Thank you so much @Helloimjohn1234 ! God bless you! c:

OpenStudy (solomonzelman):

same